# Structure from Noise: Confirmation Bias in Particle Picking in Structural Biology

### v2

@@ Abstract
Here, we develop a mathematical framework for analyzing bias in template matching-based detection with concrete applications to cryo-EM and cryo-ET. We study this bias through two downstream tasks: (i) maximum-likelihood estimation of class means in a Gaussian mixture model (GMM) and (ii) 3D volume reconstruction from the extracted particle stack. We show that when template matching is applied to pure noise, then under broad noise models, the resulting maximum-likelihood estimates converge asymptotically to deterministic, noise-dependent transforms of the user-specified templates, yielding a structure from noise effect. We further characterize how the resulting bias depends on the noise statistics, sample size, dimension, and detection threshold. Finally, controlled experiments using standard cryo-EM software corroborate the theory, demonstrating reproducible structure from noise artifacts in low-SNR data.
@@ 1.1 Examples of confirmation bias in structural biology
This phenomenon played a key role in a significant scientific controversy over the structure of an HIV molecule [33, 24, 55, 53, 32] and has since served as a cautionary tale for methodological overfitting. For a detailed statistical treatment, see [5, 6]. While often cited in discussions of particle picking using template-matching, it is crucial to note that the Einstein from Noise effect arises from a distinct mechanism: in template-matching particle picking methods, confirmation bias arises from selecting micrograph patches that most closely match the templates (essentially a form of selection bias, which is the topic of this work). In contrast, the Einstein from Noise effect emerges from aligning already extracted particles to a reference template. We further elaborate on this distinction in Appendix E.
@@
This study presents a theoretical and empirical framework to systematically investigate confirmation bias during the crucial particle-picking stage of cryo-EM and cryo-ET. Particle picking is the process of extracting 2D particle projections from noisy micrographs (in cryo-EM, Figure 1(b–c)) or 3D sub-tomograms from noisy tomograms (in cryo-ET, Figure 2(b)). While there are many different particle-picking algorithms, most fall into two main categories: template-matching algorithms (e.g., [16, 25, 34, 13]), and deep learning algorithms (e.g., [12, 56]). Template-matching methods work by scanning the micrograph or tomogram with a predefined reference structure (or template) and selecting regions that correlate highly with it. This technique has been a foundational tool in cryo-EM and cryo-ET since its inception, and it remains a widely used approach today [39]. In contrast, deep learning methods train neural networks, often convolutional neural networks (CNNs), to detect particles based on features learned directly from annotated training data. While both approaches perform well under moderate-to-high SNR, we demonstrate that in low-SNR conditions, common to both cryo-EM and cryo-ET, they can introduce systematic structural biases.
Figure 1: The cryo-electron microscope (cryo-EM) computational pipeline and structure from noise in particle picking. (a) Schematic cryo-EM workflow from 2D micrographs to a 3D reconstruction [9]: preprocessing (motion correction and CTF estimation/correction) is followed by particle picking, 2D classification, and 3D ab initio reconstruction and refinement. Red boxes mark stages particularly vulnerable to confirmation bias. (b–c) Structure from noise under template-matching picking. The template-matching picker (Algorithm 1) is applied to pure-noise micrographs using two different 3D templates: (b) Plasmodium falciparum 80S ribosome [59] and (c) beta-galactosidase [8]. The extracted particles are then processed by 2D classification (RELION VDAM [27]), yielding 2D class centers that closely resemble the picking templates despite the noise-only input. Settings: $L=20$ template projections (and $L=20$ 2D classification classes), patch size $d=48\times 48$, $M=2\times 10^{5}$ extracted particles (from 1000 micrographs), and threshold $T=3.8$; see Section 5 for full parameters.
@@ 1.2 Main results and work structure
To systematically analyze confirmation bias at the particle-picking stage, we study both classical template matching and deep learning-based particle pickers. To quantify the resulting bias, we introduce a metric that measures the similarity between the extracted particles and the original templates, defined via the outputs of two key downstream tasks: maximum-likelihood estimation of Gaussian mixture model (GMM) class centers (i.e., the statistical model governs the 2D classification task) and 3D volume reconstruction. We show that, when picking is applied to pure noise (under several models, including white, spherically symmetric, and stationary Gaussian models), the resulting maximum-likelihood estimates converge to deterministic transforms of the picking templates. This reveals a form of model bias: even in the absence of any true particles, the reconstruction is systematically shaped by the user-specified templates, rather than by genuine structure in the data.
@@ 4 Applications to cryo-EM and cryo-ET
The 2D classification step in cryo-EM is commonly modeled using a GMM, exactly of the form introduced in (2.10). Consequently, the results obtained for GMM-based class averaging developed in Section 3 apply directly to the cryo-EM 2D classification pipeline. We note, however, that the GMM model (2.10) omits a standard component of practical cryo-EM workflows: in-plane alignment, whereby each particle is rotated and translated in the image plane to best match a class average [27]. We exclude in-plane alignment from the theoretical model to preserve analytical tractability. However, as shown in Section 5, empirical experiments indicate that the core confirmation bias phenomena identified under the simplified model (2.10) persist even when in-plane alignment is included.
@@ 5.1 Particle picking based on template-matching
The theoretical model in Section 3 does not account for in-plane alignment and assumes that the number of templates used during particle picking matches the number of 2D classes, an assumption that is not expected to be met in practice. In this section, we consider a more realistic cryo-EM pipeline that includes in-plane alignment and allows the numbers of templates and 2D classes to differ.
@@
We also performed ab initio 3D reconstruction using RELION’s VDAM algorithm. In this context, ab initio reconstruction refers to generating an initial 3D structure without relying on any external reference, thereby reducing potential sources of bias. We employed RELION’s VDAM algorithm for this task with the following parameters: $L=1$ class, 200 iterations, a regularization parameter of 4, no CTF correction, and no symmetry constraints (i.e., $C1$ symmetry).
@@ Results.
Figure 5 presents results from the empirical setup in which the micrograph contains genuine particle projections, but the template used for particle picking corresponds to a different structure. In this scenario, the 2D class averages reflect a mixture of two competing sources: the true underlying particles and the mismatched template. As a result, while some class averages recover features of the actual particles, others resemble the template used for picking. Similarly, the ab initio 3D reconstruction derived from these particles exhibits structural elements from both the true volume and the template. This highlights the risk of introducing structural bias at the particle-picking stage, where even in the presence of a signal, the use of an incorrect template can propagate artifacts into downstream analyses and compromise the reliability of the final reconstruction.
Figure 5: Confirmation bias in particle picking when particles are present. Micrographs contain true particle projections of beta-galactosidase [8] corrupted by additive white Gaussian noise (SNR $\approx 1/25$), and particle picking is performed by template matching using either: (a) projections of the true beta-galactosidase structure, or (b) projections of an unrelated structure, the Plasmodium falciparum 80S ribosome [59]. In both cases, the extracted particles are processed by 2D classification and ab initio 3D reconstruction in RELION (VDAM) [27]. With matched templates (a), the 2D classes and 3D map resemble the ground truth (PCC =0.51=0.51, Fourier Shell Correlation (FSC) resolution =12.84=12.84 pixels); with mismatched templates (b), spurious ribosome-like classes appear and the 3D reconstruction becomes a hybrid, with degraded agreement (PCC =0.33=0.33, FSC resolution =27.55=27.55 pixels). Settings: patch size $d=36\times 36$, $M=2\times 10^{5}$ extracted particles, threshold $T=4$; VDAM ran for 200 iterations without CTF correction.
@@ 5.2 Particle picking based on Topaz neural network
Deep learning–based particle pickers are now widely used in cryo-EM and cryo-ET, with Topaz being a prominent example [12]. Unlike template matching, which introduces bias through more intricate mechanisms, such as explicit similarity to predefined structures, deep neural networks can inherit bias from their training data, architecture, and optimization dynamics. As a result, they may still exhibit confirmation bias by preferentially selecting particles that resemble features emphasized during training. To probe this effect, we use the pre-trained Topaz model [12] (without additional training) on micrographs of size $2048\times 2048$, down-sampled by a factor $s=2$ inside Topaz. During extraction we set the particle radius to $r=9$ (corresponding to $d=36\times 36$ in the original grid) and a coordinate scaling factor of $x=2$ to account for the down-sampling.
@@
Figure 6(a-c) shows that Topaz also induces structured bias. The mean of the selected particles by Topaz in Figure 6(a) displays a faint central circular feature, whereas under an unbiased model the mean of pure-noise patches should converge to zero. A similar central pattern appears in Figure 6(b), which shows the mean of particles selected by template matching with a large, heterogeneous template set, suggesting that both methods imprint structure on the picks. In contrast, the mean of uniformly random patches from the same noise micrographs (Figure 6(c)) is essentially zero and serves as an unbiased reference.
@@
To assess the impact on 3D refinement, we perform a controlled experiment in which pure-noise micrographs are split into two independent halves. Each half is processed with one of three picking strategies: template matching, Topaz, or random selection. Independent ab initio 3D reconstructions are then computed from each half using RELION’s VDAM algorithm, and their consistency is measured by the Fourier Shell Correlation (FSC), shown in Figure 6(d). Both Topaz and template matching yield high FSC curves between the halves, indicating that each method produces reproducible, method-dependent structure from noise, whereas random selection yields near-zero FSC, as expected under an unbiased model. Among the three, template matching exhibits the strongest bias, followed by Topaz, with random picking providing a near-unbiased baseline. These results further demonstrate that the FSC gold standard can be misled by systematic errors, in line with the observations of [52].
@@
Figure 6: Confirmation bias in Topaz particle picking. (a–c) Mean of picked particles from pure-noise micrographs: (a) Topaz picks yield a nonzero mean with a prominent central circular feature; (b) template matching with a large bank of Plasmodium falciparum 80S ribosome projection templates [59] produces a similar pattern; (c) random picks average to zero. (d) Bias assessment via half-map reproducibility: noise micrographs are split into two independent halves; on each half we pick $M=10^{5}$ particles using Topaz, template matching, or random selection, and reconstruct two half-maps using RELION VDAM. We report the Fourier Shell Correlation (FSC) between the half-maps. Settings: patch size $d=36\times 36$; VDAM ran for 200 iterations without CTF correction.
@@ 6 Discussion and Outlook
To contextualize the confirmation bias analyzed in this work, we compare it to other sources of bias that arise throughout the cryo-EM and cryo-ET reconstruction pipeline as elaborated in Appendix E. Here, we discuss strategies for mitigating this bias.
Previous studies have explored the detection and mitigation of various forms of bias in cryo-EM. For example, recent work [52] has proposed several practical strategies to reduce reconstruction bias, including identifying bias at the level of estimated parameters, performing multiple independent reconstructions, averaging parameter estimates to reduce variance, applying phase randomization at the image level, and validating algorithms across diverse preprocessing and reconstruction settings. Despite these advances, the particle-picking stage remains a particularly vulnerable and relatively understudied source of bias. Here, we focus specifically on computational strategies to mitigate bias during particle picking, incorporating both practical considerations and new theoretically grounded approaches.

### v3

@@ Abstract
Here, we develop a mathematical framework for analyzing bias in template matching-based detection with concrete applications to cryo-EM and cryo-ET. We study this bias through two downstream tasks: (i) maximum-likelihood estimation of class means in a Gaussian mixture model and (ii) 3D volume reconstruction from the extracted particle stack. We show that when template matching is applied to pure noise, then under broad noise models, the resulting maximum-likelihood estimates converge asymptotically to deterministic, noise-dependent transforms of the user-specified templates, yielding a structure from noise effect. We further characterize how the resulting bias depends on the noise statistics, sample size, dimension, and detection threshold. Finally, controlled experiments using standard cryo-EM software corroborate the theory, demonstrating reproducible structure from noise artifacts in low-SNR data.
@@ 1.1 Examples of confirmation bias in structural biology
This phenomenon played a key role in a significant scientific controversy over the structure of an HIV molecule [32, 22, 53, 51, 31] and has since served as a cautionary tale for methodological overfitting. For a detailed statistical treatment, see [5, 6]. While often cited in discussions of particle picking using template-matching, it is crucial to note that the Einstein from Noise effect arises from a distinct mechanism: in template-matching particle picking methods, confirmation bias arises from selecting micrograph patches that most closely match the templates (essentially a form of selection bias, which is the topic of this work). In contrast, the Einstein from Noise effect emerges from aligning already extracted particles to a reference template. We further elaborate on this distinction in Appendix 6.1.
@@
This study presents a theoretical and empirical framework to systematically investigate confirmation bias during the crucial particle-picking stage of cryo-EM and cryo-ET. Particle picking is the process of extracting 2D particle projections from noisy micrographs (in cryo-EM, Figure 1(b–c)) or 3D subtomograms from noisy tomograms (in cryo-ET, Figure 2(b)). While there are many different particle-picking algorithms, most fall into three broad categories: template-matching algorithms (e.g., [16, 23, 33, 13]), template-free methods based on generic shape cues such as blob- or Laplacian-of-Gaussian (LoG)-based picking, and deep learning algorithms (e.g., [12, 54]); see Section 6 for further discussion.
Template-matching methods work by scanning the micrograph or tomogram with a predefined reference structure (or template) and selecting regions that correlate highly with it. Template-free methods, by contrast, avoid explicit structural references and instead rely on coarse geometric or intensity-based cues to identify candidate particles. Deep learning methods train neural networks, often convolutional neural networks (CNNs), to detect particles based on features learned directly from annotated training data. While these approaches often perform well under moderate-to-high SNR, we show that in low-SNR conditions, common to both cryo-EM and cryo-ET, they can introduce systematic structural biases.
Figure 1: The cryo-electron microscope (cryo-EM) computational pipeline and structure from noise in particle picking. (a) Schematic cryo-EM workflow from 2D micrographs to a 3D reconstruction [9]: preprocessing (motion correction and CTF estimation/correction) is followed by particle picking, 2D classification, and 3D ab initio reconstruction and refinement. Red boxes mark stages particularly vulnerable to confirmation bias. (b–c) Structure from noise under template-matching picking. The template-matching picker (Algorithm 1) is applied to pure-noise micrographs using two different 3D templates: (b) Plasmodium falciparum 80S ribosome [57] and (c) beta-galactosidase [8]. The extracted particles are then processed by 2D classification (RELION VDAM [25], without in-plane alignment, yielding 2D class centers that closely resemble the picking templates despite the noise-only input. Settings: $L=20$ template projections (and $L=20$ 2D classification classes), patch size $d=48\times 48$, $M=2\times 10^{5}$ extracted particles (from 1000 micrographs), and threshold $T=3.8$; ; see Section 5 for an analogous experiment with in-plane alignment, together with the full set of parameters.
@@ 1.2 Main results and work structure
While prior work [4, 5, 6] studied confirmation bias arising in downstream classification, where the bias is induced at the level of the estimation and classification procedure itself, the present work focuses on a different statistical mechanism: confirmation bias introduced already at the particle-picking stage, where template matching selects noise patches that are subsequently passed to downstream reconstruction. Thus, the main novelty of this work is the analysis of particle-picking-induced selection bias and how it affects downstream tasks. Section 6.1 and Figure 8 further compare this mechanism with other forms of confirmation bias arising elsewhere in the cryo-EM/cryo-ET pipeline.
To systematically analyze confirmation bias at the particle-picking stage, we study both classical template matching and deep learning-based particle pickers. To quantify the resulting bias, we introduce a metric that measures the similarity between the extracted particles and the original templates, defined via the outputs of two key downstream tasks: maximum-likelihood estimation of Gaussian mixture model (GMM) class centers (i.e., the statistical model that governs the 2D classification task) and 3D volume reconstruction. We show that, when picking is applied to pure noise (under several models, including white, spherically symmetric, and stationary Gaussian models), the resulting maximum-likelihood estimates converge to deterministic transforms of the picking templates. This reveals a form of model bias: even in the absence of any true particles, the reconstruction is systematically shaped by the user-specified templates, rather than by genuine structure in the data.
@@ 2.3 Probabilistic models for the input patches
Informally, this condition provides a mathematical model for short-range dependence: local regions of a signal (e.g., a micrograph or tomogram) are expected to be correlated, but the dependence weakens with distance, so sufficiently separated patches behave almost independently. Assumption (2.7) strengthens this by requiring that the dependence decay fast enough for empirical averages over many selected patches to exhibit the same qualitative concentration behavior as in the independent case. Such assumptions are standard in the analysis of stochastic processes and are natural here for analyzing the spatially correlated noise model with limited range.
@@ 4 Applications to cryo-EM and cryo-ET
The 2D classification step in cryo-EM is commonly modeled using a GMM, exactly of the form introduced in (2.10). Consequently, the results obtained for GMM-based class averaging developed in Section 3 apply directly to the cryo-EM 2D classification pipeline. We note, however, that the simplified GMM model (2.10) omits two standard aspects of practical cryo-EM workflows. First, in-plane alignment is typically performed, whereby each particle is rotated and translated in the image plane to best match a class average [25]. Second, the number of downstream 2D classes need not equal the number of picking templates. In the theory, we exclude in-plane alignment and take $K=L$ to isolate the particle-picking-induced bias in its clearest analytically tractable form. These simplifications are not essential to the qualitative mechanism: when $K\neq L$, template-based selection still induces a biased conditional distribution on the extracted patches, but the downstream GMM no longer yields a one-to-one correspondence between templates and fitted classes; if $K<L$, several template-induced modes may merge, whereas if $K>L$, a single such mode may split into multiple fitted components. As shown empirically in Section 5, the core confirmation-bias phenomena persist in these more realistic settings.
@@ 5.1 Particle picking based on template-matching
The theoretical model in Section 3, excludes in-plane alignment and assumes that the number of templates used during particle picking matches the number of 2D classes, an assumption that is generally not satisfied in practice. In this section, we consider a more realistic cryo-EM pipeline that includes in-plane alignment and allows the numbers of templates and 2D classes to differ.
@@
We also performed ab initio 3D reconstruction using RELION’s VDAM algorithm. In this context, ab initio reconstruction refers to generating an initial 3D structure without relying on any external reference, thereby reducing potential sources of bias. We employed RELION’s VDAM algorithm for this task with the following parameters: $L=1$ class, 200 iterations, a regularization parameter of 4, no CTF correction, and no symmetry constraints (i.e., $C1$ symmetry). Unlike the idealized setting of Corollary 4.2, this empirical reconstruction does not use the same discrete rotation grid as the one used during template matching. Rather, the experiment is intentionally carried out in a more realistic regime, in which the picking and reconstruction stages use different angular discretizations. This allows us to assess the extent to which the predicted confirmation-bias phenomenon persists under practical grid mismatch.
@@ Results.
Figure 5 presents results from the empirical setup in which the micrograph contains genuine particle projections, but the template used for particle picking corresponds to a different structure. In this scenario, the 2D class averages reflect a mixture of two competing sources: the true underlying particles and the mismatched template. As a result, while some class averages recover features of the actual particles, others resemble the template used for picking. Similarly, the ab initio 3D reconstruction derived from these particles exhibits structural elements from both the true volume and the template. This highlights the risk of introducing structural bias at the particle-picking stage, where even in the presence of a signal, the use of an incorrect template can propagate artifacts into downstream analyses and compromise the reliability of the final reconstruction. In addition, we note that the use of a markedly dissimilar template in Figure 5 is intended to make the bias mechanism visually transparent. In practice, however, the more dangerous regime may be one in which the incorrect template is structurally similar to the true particle, since the resulting bias can be substantially harder to detect.
Figure 5: Confirmation bias in particle picking when particles are present. Micrographs contain true particle projections of beta-galactosidase [8] corrupted by additive white Gaussian noise (SNR $\approx 1/25$), and particle picking is performed by template matching using either: (a) projections of the true beta-galactosidase structure, or (b) projections of an unrelated structure, the Plasmodium falciparum 80S ribosome [57]. In both cases, the extracted particles are processed by 2D classification and ab initio 3D reconstruction in RELION (VDAM) [25]. With matched templates (a), the 2D classes and 3D map resemble the ground truth (PCC =0.69=0.69, Fourier Shell Correlation (FSC) resolution =2.99=2.99 pixels); with mismatched templates (b), spurious ribosome-like classes appear and the 3D reconstruction becomes a hybrid, with degraded agreement (PCC =0.59=0.59, FSC resolution =4.65=4.65 pixels). Settings: patch size $d=36\times 36$, $M=2\times 10^{5}$ extracted particles, threshold $T=4$; VDAM ran for 200 iterations without CTF correction.
@@ 5.2 Particle picking based on Topaz neural network
Deep learning–based particle pickers are now widely used in cryo-EM and cryo-ET, with Topaz being a prominent example [12]. Unlike template matching, which introduces bias through more intricate mechanisms, such as explicit similarity to predefined structures, deep neural networks can inherit bias from their training data, architecture, and optimization dynamics. As a result, they may still exhibit confirmation bias by preferentially selecting particles that resemble features emphasized during training. To probe this effect, we consider both the pre-trained Topaz model [12], used without additional training, and Topaz models trained on user-specified particle sets.
Pre-trained Topaz network.
For the pre-trained Topaz network, we applied Topaz to micrographs of size $2048\times 2048$, using an internal down-sampling factor of $s=2$. During extraction, we set the particle radius to $r=9$, corresponding to a patch size of $d=36\times 36$ on the original grid, and used a coordinate scaling factor of $x=2$ to account for the down-sampling. We used a score threshold of 1-1 for the Topaz output, corresponding to a predicted particle probability of $p\geq 0.25$. Although the default threshold in Topaz and RELION is typically 0 (that is, $p\geq 0.5$), in practice users often lower this threshold to improve recall, especially for challenging low-SNR datasets. We adopted this more permissive setting to extract a sufficiently large particle stack ($M=2\times 10^{5}$) within a reasonable runtime, enabling a direct comparison with the template-matching experiments in both dataset size and effective SNR. Thus, the observed structure-from-noise artifacts arise under realistic conditions, rather than from forcing the network to select arbitrary low-confidence patches.
@@
Figure 6(a-c) shows that the pre-trained Topaz network induces structured bias. The mean of the selected particles by Topaz in Figure 6(a) displays a faint central circular feature, whereas under an unbiased model the mean of pure-noise patches should converge to zero. A similar central pattern appears in Figure 6(b), which shows the mean of particles selected by template matching with a large, heterogeneous template set, suggesting that both methods imprint structure on the picks. In contrast, the mean of uniformly random patches from the same noise micrographs (Figure 6(c)) is essentially zero.
@@
To assess the impact on 3D refinement, we perform a controlled experiment in which pure-noise micrographs are split into two independent halves. Each half is processed with one of three picking strategies: template matching, Topaz, or random selection. Independent ab initio 3D reconstructions are then computed from each half using RELION’s VDAM algorithm, and their consistency is measured by the Fourier Shell Correlation (FSC), shown in Figure 6(d). Both Topaz and template matching yield high FSC curves between the halves, indicating that each method produces reproducible, method-dependent structure from noise. In contrast, random selection yields near-zero FSC, as expected under an unbiased model. Among the three, template matching exhibits the strongest bias, followed by Topaz. These results further demonstrate that the FSC gold standard can be misled by systematic errors, in line with the observations of [50].
@@
Figure 6: Confirmation bias in pre-trained Topaz particle picking. (a–c) Mean of picked particles from pure-noise micrographs: (a) Topaz picks yield a nonzero mean with a prominent central circular feature; (b) template matching with a large bank of Plasmodium falciparum 80S ribosome projection templates [57] produces a similar pattern; (c) random picks average to zero. (d) Bias assessment via half-map reproducibility: noise micrographs are split into two independent halves; on each half we pick $M=10^{5}$ particles using Topaz, template matching, or random selection, and reconstruct two half-maps using RELION VDAM. We report the Fourier Shell Correlation (FSC) between the half-maps. Settings: patch size $d=36\times 36$; VDAM ran for 200 iterations without CTF correction.
Figure 7: Confirmation bias in trained Topaz particle picking when true particles are present. Micrographs contain true particle projections of beta-galactosidase [8] corrupted by additive white Gaussian noise. Particle picking is performed using a Topaz model trained on either: (a) projections of the true beta-galactosidase structure (matched training data), or (b) projections of an unrelated structure, the Plasmodium falciparum 80S ribosome [57] (mismatched training data). In both cases, the extracted particles are processed by 2D classification and ab initio 3D reconstruction in RELION (VDAM) [25]. With the matched training data (a), the 2D classes and 3D map successfully recover the beta-galactosidase structure (PCC =0.24=0.24, FSC resolution =6.84=6.84 pixels). Conversely, when trained on the mismatched structure (b), the network preferentially selects noise features that align with its learned prior. As a result, the 3D reconstruction of the true underlying signal is significantly degraded.
Trained Topaz network.
To compare deep-learning-based picking with classical template matching under realistic conditions, we tested whether Topaz’s learned prior can override weak true signal when real particles are present. Following the setup of Figure 5, we generated synthetic micrographs containing true beta-galactosidase projections at low SNR ($\approx 1/25$). Starting from the same vanilla pre-trained Topaz model used in the previous experiment, we re-trained two networks: one on positive examples from beta-galactosidase, and one on an unrelated structure, the Plasmodium falciparum 80S ribosome. In both cases, training was run for 10 epochs on 20 synthetic micrographs, each containing 200 particle projections embedded in white Gaussian noise at the same low-SNR level as the evaluation data.
As shown in Figure 7, when Topaz is trained on the mismatched ribosome data, the extracted particles yield 2D class averages and a 3D reconstruction with clear ribosome-like features. The reconstruction quality is also substantially worse than in the matched-training case (PCC 0.08 versus 0.24, and FSC resolution 17.08 pixels versus 6.84 pixels). This shows that, much like template matching, a dedicated neural-network picker can be strongly affected by the specific structural prior imposed during training. Rather than introducing only a generic bias toward centered, particle-like blobs, the network preferentially selects low-SNR data that resemble its training distribution, thereby propagating structure-specific hallucinations into the downstream pipeline.
@@ 6 Discussion and Outlook
To place the bias analyzed in this work in context, we compare it with other sources of bias that arise throughout the cryo-EM and cryo-ET reconstruction pipelines. We then discuss mitigation strategies and conclude with several promising directions for future research.
6.1 Different forms of confirmation bias in cryo-EM
Several stages of the cryo-EM and cryo-ET pipeline are vulnerable to confirmation bias [50, Section 4.1], including particle picking, 2D classification, and 3D reconstruction. The present work focuses on particle picking [44, 21, 12, 20], while related biases in low-SNR EM-based reconstruction have been analyzed in [6]. These stages introduce bias through a different statistical mechanism. Figure 8 compares confirmation bias in particle picking with that in 2D classification. Below, we contrast the Einstein from Noise phenomenon from Section 1 and the bias observed in 2D classification [4] with the bias arising from particle picking.
The Einstein from Noise phenomenon.
Although frequently mentioned in the context of template-based particle picking, the Einstein from Noise effect arises from a distinct model. In the Einstein from Noise experiment, all observations are aligned and averaged with respect to a fixed template. In contrast, particle picking selects a subset of patches with the highest template correlations. As a result, particle picking produces reconstructions that asymptotically match the template up to global scaling (Theorems 3.1 and 4.2), whereas Einstein from Noise yields an average that is only positively correlated with the template but not identical [5].
In later stages of cryo-EM processing, such as 2D classification and 3D volume estimation, iterative algorithms often use predefined templates for initialization. When applied to data with no true signal, these methods may still produce output that resembles the initial templates, even though the input data provides no evidence for such structure [4].
Figure 8(b) illustrates this effect in 2D classification. Here, the input consisted entirely of pure noise, and the initial templates were $L=30$ distinct projections of the Beta-galactosidase structure of size $d=36\times 36$. A total of $M=5\times 10^{5}$ particles were processed by EM-based 2D classification for 20 iterations. Even without signal, the resulting class averages remain positively correlated with the templates. Unlike template-based particle picking, however, they do not reproduce the templates exactly, although they remain visually similar. Many modern 2D classification and 3D ab initio reconstruction methods use reference-free initialization, such as random initialization, to mitigate this bias [25]. Nevertheless, similar confirmation effects can still arise when templates are introduced explicitly or implicitly.
6.2 Strategies for mitigating confirmation bias
Previous studies have explored the detection and mitigation of various forms of bias in cryo-EM. For example, recent work [50] has proposed several practical strategies to reduce reconstruction bias, including identifying bias at the level of estimated parameters, performing multiple independent reconstructions, averaging parameter estimates to reduce variance, applying phase randomization at the image level, and validating algorithms across diverse preprocessing and reconstruction settings. Similarly, an additional work [30] provided an empirical analysis of template bias in high-resolution 2D template matching and proposed practical conditions and diagnostics to reduce template-induced bias in reconstruction. Despite these advances, the particle-picking stage remains particularly vulnerable. Here, we focus specifically on computational strategies to mitigate bias during particle picking, incorporating both practical considerations and new theoretically grounded approaches.